OpenStudy (anonymous):

2 points Save Find the rate of change between the two points: (4, 10) and (6, 15) where x is in minutes and y is in miles. 4 miles per minute 4 minutes per mile 2.5 miles per minute 2.5 minutes per mile

OpenStudy (inkyvoyd):

SLOPE

OpenStudy (inkyvoyd):

Whoopes. Deltay/delta x

OpenStudy (anonymous):

(4, 10) and (6, 15) \[\frac{15-10}{6-4}=\frac 5 2\]=2.5 milles per minutes

OpenStudy (inkyvoyd):

Find the slope of the line, and that's your answer.